OpenStudy (anonymous):

A bag contains 4 red marbles and 5 blue marbles. What is the probability of randomly selecting a blue marble and then, without replacing it, selecting a red marble?

OpenStudy (anonymous):

5/9 x 4/8 = 5/9 x 1/2 = 5/18 Notice how the denominator in the second factor is 8. That's because there are only 8 marbles left at that point, of which there are still 4 red ones.

OpenStudy (anonymous):

where did you get 1/2?

OpenStudy (anonymous):

oh wait 1/2 is 4/8 simplified, right?

OpenStudy (anonymous):

yes.